Prove or disprove: Every subgroup of the integers has finite index.


1
Expert's answer
2022-02-27T16:33:27-0500

Let "H" be a subgroup of "\mathbb{Z}". First of all, we study apart the case of "H=\{0\}". In this case "H" has an infinite index, but this is a degenerate case.

Now let "H" be a non-trivial subgroup, then "H" has a non-zero element "x" with a minimal absolute value "|x|>0". Let us prove that "H=x\mathbb{Z}". First, "x\mathbb{Z}\subseteq H". Second, suppose that there is "y\in H" such that "y\neq a\cdot x" with "a\in \mathbb{Z}". The euclidian division of "y" by "x" gives us a remainder "r\in H" (as "r=y-a\cdot x" and both terms are in "H"), but by the properties of a euclidian division "|r|<|x| \text{ and } \; r\neq 0" as "y\neq a\cdot x". This condraticts the fact that "x\in H" is of minimal absolute value. Therefore, "H=x\cdot \mathbb{Z}" and we easily deduce that the index of "H" is "|x|" and so is finite.

We conclude that the index of every non-trivial subgroup of the integers is finite. For the trivial subgroup this is obviously false.